OpenStudy (anonymous):

Solve for X. -x/7 - 9 = -16 Why is the answer 49 and not -103?

OpenStudy (anonymous):

why do u think that the answer should be -103??????

OpenStudy (anonymous):

because if times -16 * 7 that would be -112 then + 9 = -103

OpenStudy (shamil98):

\[\huge -\frac{ x }{ 7 } - 9 = -16\] you add 9 to both sides first. then multiply by -7 to both sides.. \[\huge -\frac{ x }{ 7 } = -7\] \[\huge \cancel{-7} * \frac{ x }{ \cancel{-7} } = -7 * -7\] \[\huge x= 49\]

OpenStudy (anonymous):

is that a rule to add first then multiply?

OpenStudy (anonymous):

do u know how to add integers with fractions??????

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then first calculate this only